A financial planner wants to invest $8000, some in stock earning 15% annually and the rest in bonds earning 6% annually. How much should be invested at each rate to get a return of $930 annually from the two investments?
Answer by checkley77(12844)

.15x+.06(8,000-x)=930
.15x+480-.06x=930
.09x=930-480
.09x=450
x=450/.09
x=5,000 invested @ 15%.
8,000-5,000=3,000 invested @ 6%
Proof:
.15*5,000+.06*3,000=930
750+180=930
930=930